OpenStudy (anonymous):

Find the slopes of the asymptotes of a hyperbola with the following equation

OpenStudy (anonymous):

y^2/36-x^2/4=1

OpenStudy (anonymous):

X^2/4

OpenStudy (anonymous):

and its minus not multiply

OpenStudy (jhannybean):

ohh nvm.

OpenStudy (anonymous):

hyperbola right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

form is \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\] so you know it looks like this